# Diffusion-based Extreme Image Compression with Compressed Feature Initialization

### v1

@@ Abstract
Diffusion-based extreme image compression methods have achieved impressive performance at extremely low bitrates. However, constrained by the iterative denoising process that starts from pure noise, these methods are limited in both fidelity and efficiency. To address these two issues, we present Relay Residual Diffusion Extreme Image Compression (RDEIC), which leverages compressed feature initialization and residual diffusion. Specifically, we first use the compressed latent features of the image with added noise, instead of pure noise, as the starting point to eliminate the unnecessary initial stages of the denoising process. Second, we design a novel relay residual diffusion that reconstructs the raw image by iteratively removing the added noise and the residual between the compressed and target latent features. Notably, our relay residual diffusion network seamlessly integrates pre-trained stable diffusion to leverage its robust generative capability for high-quality reconstruction. Third, we propose a fixed-step fine-tuning strategy to eliminate the discrepancy between the training and inference phases, further improving the reconstruction quality. Extensive experiments demonstrate that the proposed RDEIC achieves state-of-the-art visual quality and outperforms existing diffusion-based extreme image compression methods in both fidelity and efficiency. The source code will be provided in https://github.com/huai-chang/RDEIC.
1 Introduction
@@
Extreme image compression is becoming increasingly important with the growing demand for efficient storage and transmission of images where storage capacity or bandwidth is limited, such as in satellite communications and mobile devices. Conventional compression standards like JPEG (Wallace, 1991), BPG (Bellard, 2014) and VVC (Bross et al., 2021) rely on hand-crafted rules and block-based redundancy removal techniques, leading to severe blurring and blocking artifacts at low bitrates. Hence, there is an urgent need to explore extreme image compression methods.
@@
In recent years, learned image compression methods have attracted significant interest, outperforming conventional codecs. However, distortion-oriented learned compression methods (Xie et al., 2021; Zhu et al., 2021; Liu et al., 2023; Li et al., 2024a) optimize for the rate-distortion function alone, resulting in unrealistic reconstructions at low bitrates, typically manifested as blurring or over-smoothing. Perceptual-oriented learned compression methods (Agustsson et al., 2019; Mentzer et al., 2020; Muckley et al., 2023; Yang & Mandt, 2023) introduce generative models, such as generative adversarial networks (GANs) (Goodfellow et al., 2014) and diffusion models (Ho et al., 2020), to enhance the perceptual quality of reconstructions. However, these methods are optimized for medium to high bitrates instead of extremely low bitrates such as below 0.1 bpp. As a result, these methods experience significant quality degradation when the compression ratio is increased.
Figure 1: Qualitative comparison between the proposed RDEIC and recent state-of-the-art methods: MS-ILLM (Muckley et al., 2023), Text+Sketch (Lei et al., 2023), PerCo (Careil et al., 2024), and DiffEIC (Li et al., 2024b). For diffusion-based methods, the number of denoising steps is written after the name, e.g. PerCo-20 and Ours-5 means 20 diffusion steps are used by PerCo and 5 diffusion steps are used by our method.
@@
Recently, diffusion-based extreme image compression methods (Lei et al., 2023; Careil et al., 2024; Li et al., 2024b) leverage the robust generative ability of pre-trained text-to-image (T2I) diffusion models, achieving superior visual quality at extremely low bitrates. Nonetheless, these methods are constrained by the inherent characteristics of diffusion models. Firstly, these methods rely on an iterative denoising process to reconstruct raw images from pure noise which is inefficient for inference (Li et al., 2024b). Secondly, initiating the denoising process from pure noise introduces significant randomness, compromising the fidelity of the reconstructions (Careil et al., 2024). Thirdly, there is a discrepancy between the training and inference phases. During training, each time-step is trained independently, which is well-suited for image generation tasks where diversity (or randomness) is encouraged (Ho et al., 2020). However, this training approach is not optimal for image compression where consistency between the reconstruction and the raw image is crucial.
@@
In this work, we propose Relay Residual Diffusion Extreme Image Compression (RDEIC) to overcome the three limitations mentioned above. To overcome the first two limitations, inspired by Relay Diffusion (Teng et al., 2024), we use the compressed latent features with added noise as the starting point, which significantly reduces the number of denoising steps required for reconstruction and enhances the reconstruction fidelity since providing a better starting point. We also propose a novel relay residual diffusion process that reconstructs the raw image by iteratively removing the added noise and the residual between the compressed and target latent features (i.e., the features of the raw image in the latent diffusion space). Furthermore, we integrate pre-trained stable diffusion into our compression framework to leverage its robust generative capability for high perceptual reconstruction. To address the third limitation, we introduce a fixed-step fine-tuning strategy to eliminate the discrepancy between the training and inference phases. By fine-tuning RDEIC throughout the entire reconstruction process, we further improve the reconstruction quality. Moreover, to meet users’ diverse requirements, we introduce a controllable detail generation method that achieves a trade-off between smoothness and sharpness by adjusting the intensity of high-frequency components in the reconstructions. As shown in Fig. 1, the proposed RDEIC achieves state-of-the-art perceptual performance at extremely low bitrates, and significantly outperforms existing diffusion-based extreme image compression methods with fewer inference steps.
In summary, our contributions are as follows:
  • We propose RDEIC, a novel diffusion model for extreme image compression that accelerates the denoising process through compressed feature initialization i.e. the compressed latent features with added noise, instead of pure noise, are used as the starting point of reverse diffusion.
  • We propose a relay residual diffusion process that seamlessly integrates pre-trained stable diffusion into our compression network and removes the residual between the compressed and target latent features.
@@
  • To eliminate the discrepancy between the training and inference phases, we design a fixed-step fine-tuning strategy that refines the model through the entire reconstruction process, further improving reconstruction quality.
2 Related Work
@@
Learned Image Compression. As a pioneer work, Ballé et al. (2017) proposed an end-to-end image compression framework to jointly optimize the rate-distortion performance. Ballé et al. (2018) later introduced a hyperprior to reduce spatial dependencies in the latent representation, greatly enhancing performance. Subsequent works further improved compression models by developing various nonlinear transforms (Xie et al., 2021; He et al., 2022; Liu et al., 2023; Li et al., 2024a) and entropy models (Minnen et al., 2018; Minnen & Singh, 2020; He et al., 2021; Qian et al., 2021). However, optimization for rate-distortion alone often results in unrealistic reconstructions at low bitrates, typically manifested as blurring or over-smoothness (Blau & Michaeli, 2019). To improve perceptual quality, generative models have been integrated into compression methods. Agustsson et al. (2019) added an adversarial loss for lost details generation. Mentzer et al. (2020) explored the generator and discriminator architectures, as well as training strategies for perceptual image compression. Muckley et al. (2023) introduced a local adversarial discriminator to enhance statistical fidelity. With the advancement of diffusion models, some efforts have been made to apply diffusion models to image compression. For instance, Yang & Mandt (2023) innovatively introduced a conditional diffusion model as decoder for image compression. Kuang et al. (2024) proposed a consistency guidance architecture to guide the diffusion model in stably reconstructing high-quality images.
@@
Extreme Image Compression. In recent years, extreme image compression has garnered increasing attention, aiming to compress image to extremely low bitrates, often below 0.1 bpp, while maintaining visually acceptable image quality. Gao et al. (2023) leveraged the information-lossless property of invertible neural networks to mitigate the significant information loss in extreme image compression. Jiang et al. (2023) treated text descriptions as prior to ensure semantic consistency between the reconstructions and the raw images. Wei et al. (2024) achieved extreme image compression by rescaling images using extreme scaling factors. Lu et al. (2024) combined continuous and codebook-based discrete features to reconstruct high-quality images at extremely low bitrates. Inspired by the great success of T2I diffusion models in various image restoration tasks (Lin et al., 2023; Wang et al., 2024), some methods have incorporated T2I diffusion models into extreme image compression frameworks. Lei et al. (2023) utilized a pre-trained ControlNet (Zhang et al., 2023) to reconstruct images based on corresponding short text prompts and binary contour sketches. Careil et al. (2024) conditioned iterative diffusion models on vector-quantized latent image representations and textual image descriptions. Li et al. (2024b) combined compressive VAEs with pre-trained T2I diffusion models to achieve realistic reconstructions at extremely low bitrates. However, constrained by the inherent characteristics of diffusion models, these diffusion-based extreme image compression methods are limited in both fidelity and efficiency. In this paper, we propose a solution to these limitations through a relay residual diffusion framework and a fixed-step fine-tuning strategy.
@@
Relay Diffusion. Conventional diffusion models, such as denoising diffusion probabilistic models (DDPM) (Ho et al., 2020) and its variants, operate by progressively adding and removing Gaussian noise through a multi-step Markov process. These models have achieved remarkable results in low-resolution scenarios but face substantial challenges in terms of computational efficiency and performance when applied to higher resolutions. To overcome this, cascaded diffusion methods (Ho et al., 2022; Saharia et al., 2022) are introduced, which decompose the image generation into multiple stages, with each stage responsible for super-resolution conditioning on the previous one. However, these methods still require complete resampling at each stage, leading to inefficiencies and potential mismatches among different resolutions. Relay diffusion, as proposed by Teng et al. (2024), extends the cascaded framework by continuing the diffusion process directly from the low-resolution output rather than restarting from pure noise, which allows the higher-resolution stages to correct artifacts from earlier stages. This design is particularly well-suited for extreme image compression tasks where compressed latent features of the raw image are available. Therefore, we propose a novel relay residual diffusion process that integrates relay diffusion with diffusion-based extreme image compression method, greatly enhancing reconstruction efficiency and fidelity at extremely low bitrates.
@@
Figure 2: The proposed RDEIC. We first map a raw image $\bm{x}$ into the latent space using the encoder $\mathcal{E}$ and then perform end-to-end lossy compression to get compressed latent features $\bm{z}_{c}$. We then use $\bm{z}_{c}$ with added noise as the starting point and apply a denoising process to reconstruct the noise-free latent feature $\bm{z}_{0}$. The decoder $\mathcal{D}$ maps $\bm{z}_{0}$ back to the pixel space, to get the reconstructed image $\hat{\bm{x}}$. (a) Vanilla diffusion framework that starts from pure noise. (b) The proposed relay residual diffusion framework that starts from compressed latent features with added noise.
3 Methodology
3.1 Overall Framework
@@
Fig. 2 shows an overview of the proposed RDEIC network. We first use an encoder $\mathcal{E}$ and analysis transform $g_{a}$ to convert the input image $\bm{x}$ to its latent representation $\bm{y}$. Then we perform hyper transform coding on $\bm{y}$ with the categorical hyper model (Jia et al., 2024) and use the space-channel context model $\mathcal{C}_{m}$ to predict the entropy parameters $(\bm{\mu},\bm{\sigma})$ to estimate the distribution of quantized latent representation $\bm{\hat{y}}$ (He et al., 2022). Subsequently, the synthesis transform $g_{s}$ is used to obtain the image content dependent features $\bm{z}_{c}$. Random noise is then added to $\bm{z}_{c}$, which is the starting point for reconstructing the noise-free latent features $\bm{z}_{0}$ through an iterative denoising process. The denoising process is implemented by a frozen pre-trained noise estimator $\epsilon_{sd}$ of stable diffusion with trainable control network for intermediate feature modulation. Finally, the reconstructed image $\bm{\hat{x}}$ is decoded from $\bm{z}_{0}$ using the decoder $\mathcal{D}$.
3.2 Accelerating Denoising Process with Relay Residual Diffusion
@@
Following stable diffusion, existing diffusion-based extreme image compression methods obtain the noisy latent by adding Gaussian noise with variance $\beta_{t}\in(0,1)$ to the noise-free latent features $\bm{z}_{0}$:
$\bm{z}_{t}=\sqrt{\bar{\alpha}_{t}}\bm{z}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{t},\ t=1,2,\cdots,T,$ (1)
@@
where $\epsilon_{t}\sim\mathcal{N}(0,\bm{I})$, $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$. When $t$ is large enough, the noisy latent $\bm{z}_{t}$ is nearly a standard Gaussian distribution. In practice, $T$ is typically very large, e.g., 1000, and pure noise is set as the starting point for the reverse diffusion process. However, this approach is not optimal for the image compression task, where the compressed latent features $\bm{z}_{c}$ are available.
@@
To this end, we set the starting point to $\bm{z}_{N}=\sqrt{\bar{\alpha}_{N}}\bm{z}_{c}+\sqrt{1-\bar{\alpha}_{N}}\epsilon_{N}$, where $N\ll T$. Our relay residual diffusion is thus defined as:
$\bm{z}_{n}=\sqrt{\bar{\alpha}_{n}}(\bm{z}_{0}+\eta_{n}\bm{e})+\sqrt{1-\bar{\alpha}_{n}}\epsilon_{n},\ n=1,2,\cdots,N,$ (2)
@@
where $\bm{e}$ denotes the residual between $\bm{z}_{c}$ and $\bm{z}_{0}$, i.e., $\bm{e}=\bm{z}_{c}-\bm{z}_{0}$, and $\{\eta_{n}\}_{n=1}^{N}$ is a weight sequence that satisfies $\eta_{1}\to 0$ and $\eta_{N}=1$. Since the residual $\bm{e}$ is unavailable during inference, we refer to DDIM (Song et al., 2021) and assume that $\bm{z}_{n-1}$ is a linear combination of $\bm{z}_{n}$ and $\bm{z}_{0}$:
$\bm{z}_{n-1}=k_{n}\bm{z}_{0}+m_{n}\bm{z}_{n}+\sigma_{n}\epsilon,$ (3)
where we set $\sigma_{n}=0$ for simplicity. Combining Eq. (2) and Eq. (3), we get

### v2

# RDEIC: Accelerating Diffusion-Based Extreme Image Compression with Relay Residual Diffusion

@@ Abstract
Diffusion-based extreme image compression methods have achieved impressive performance at extremely low bitrates. However, constrained by the iterative denoising process that starts from pure noise, these methods are limited in both fidelity and efficiency. To address these two issues, we present Relay Residual Diffusion Extreme Image Compression (RDEIC), which leverages compressed feature initialization and residual diffusion. Specifically, we first use the compressed latent features of the image with added noise, instead of pure noise, as the starting point to eliminate the unnecessary initial stages of the denoising process. Second, we directly derive a novel residual diffusion equation from Stable Diffusion’s original diffusion equation that reconstructs the raw image by iteratively removing the added noise and the residual between the compressed and target latent features. In this way, we effectively combine the efficiency of residual diffusion with the powerful generative capability of Stable Diffusion. Third, we propose a fixed-step fine-tuning strategy to eliminate the discrepancy between the training and inference phases, thereby further improving the reconstruction quality. Extensive experiments demonstrate that the proposed RDEIC achieves state-of-the-art visual quality and outperforms existing diffusion-based extreme image compression methods in both fidelity and efficiency. The source code and pre-trained models will be released.
Index Terms:
Image compression, compressed latent features, residual diffusion, extremely low bitrates
I Introduction
@@
Extreme image compression is becoming increasingly important with the growing demand for efficient storage and transmission of images where storage capacity or bandwidth is limited, such as in satellite communications and mobile devices. Conventional compression standards like JPEG [1], BPG [2] and VVC [3] rely on hand-crafted rules and block-based redundancy removal techniques, leading to severe blurring and blocking artifacts at low bitrates. Hence, there is an urgent need to explore extreme image compression methods.
@@
In recent years, learned image compression methods have attracted significant interest, outperforming conventional codecs. However, distortion-oriented learned compression methods [4, 5, 6, 7] optimize for the rate-distortion function alone, resulting in unrealistic reconstructions at low bitrates, typically manifested as blurring or over-smoothing. Perceptual-oriented learned compression methods [8, 9, 10, 11] introduce generative models, such as generative adversarial networks (GANs) [12] and diffusion models [13], to enhance the perceptual quality of reconstructions. However, these methods are optimized for medium to high bitrates instead of extremely low bitrates such as below 0.1 bpp. As a result, these methods experience significant quality degradation when the compression ratio is increased.
@@
Recently, diffusion-based extreme image compression methods [14, 15, 16] leverage the robust generative ability of pre-trained text-to-image (T2I) diffusion models, achieving superior visual quality at extremely low bitrates. Nonetheless, these methods are constrained by the inherent characteristics of diffusion models. Firstly, these methods rely on an iterative denoising process to reconstruct raw images from pure noise, which is inefficient for inference [16]. Secondly, initiating the denoising process from pure noise introduces significant randomness, compromising the fidelity of the reconstructions [15]. Thirdly, there is a discrepancy between the training and inference phases. During training, each time-step is trained independently, which is well-suited for image generation tasks where diversity (or randomness) is encouraged [13]. However, this training approach is not optimal for image compression where consistency between the reconstruction and the raw image is crucial.
Original VVC MS-ILLM Text+Sketch-25 PerCo-20 DiffEIC-50 RDEIC-5 (Ours)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
[…]
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
[…]
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
[…]
Figure 1: Qualitative comparisons between the proposed RDEIC and state-of-the-art methods. The number of denoising steps is written after the name, e.g. DiffEIC-50 means 50 diffusion steps are used by DiffEIC. The bpp and DISTS of each method are shown at the bottom of each image.
@@
In this work, we propose Relay Residual Diffusion Extreme Image Compression (RDEIC) to overcome the three limitations mentioned above. To overcome the first two limitations, we proposed a novel relay residual diffusion framework. Specifically, we construct the starting point using the compressed latent features combined with slight noise, transitioning between the starting point and target latent features by shifting the residual between them. This approach significantly reduces the number of denoising steps required for reconstruction while ensures that the starting point retains most of the information from the compressed features, providing a strong foundation for subsequent detail generation. To leverage the robust generative capability of pre-trained stable diffusion for extreme image compression, we derive a novel residual diffusion equation directly from stable diffusion’s diffusion equation, rather than designing a new diffusion equation from scratch as done in [17]. To address the third limitation, we introduce a fixed-step fine-tuning strategy to eliminate the discrepancy between the training and inference phases. By fine-tuning RDEIC throughout the entire reconstruction process, we further improve the reconstruction quality. Moreover, to meet users’ diverse requirements, we introduce a controllable detail generation method that achieves a trade-off between smoothness and sharpness by adjusting the intensity of high-frequency components in the reconstructions. As shown in Fig. 1, the proposed RDEIC achieves state-of-the-art perceptual performance at extremely low bitrates, and significantly outperforms existing diffusion-based extreme image compression methods with fewer inference steps.
In summary, the main contributions of this work are as follows:
1) We propose a relay residual diffusion process that effectively combine the efficiency of residual diffusion with the powerful generative capability of stable diffusion. To the best of our knowledge, we are the first to successfully integrate stable diffusion into a residual diffusion framework.
@@
2) To eliminate the discrepancy between the training and inference phases, we design a fixed-step fine-tuning strategy that refines the model through the entire reconstruction process, further improving reconstruction quality.
3) We introduce a controllable detail generation method to balance smoothness and sharpness, allowing users to explore and customize outputs according to their personal preferences.
The remainder of this paper is organized as follows. The related works are summarized in Section II. The proposed method is described in Section III. The experiment results and analysis are presented in Section IV and Section V, respectively. Finally, we conclude our work in Section VI.
II Related Work
II-A Learned Image Compression
@@
As a pioneer work, Ballé et al. [18] proposed an end-to-end image compression framework to jointly optimize the rate-distortion performance. In [19], they later introduced a hyperprior to reduce spatial dependencies in the latent representation, greatly enhancing performance. Subsequent works further improved compression models by developing various nonlinear transforms [4, 6, 7, 20] and entropy models [21, 22, 23, 24]. However, optimization for rate-distortion alone often results in unrealistic reconstructions at low bitrates, typically manifested as blurring or over-smoothness [25].
To improve perceptual quality, generative models have been integrated into compression methods. Agustsson et al. [8] added an adversarial loss for lost details generation. Mentzer et al. [9] explored the generator and discriminator architectures, as well as training strategies for perceptual image compression. In [10], Muckley et al. introduced a local adversarial discriminator to enhance statistical fidelity. With the advancement of diffusion models, some efforts have been made to apply diffusion models to image compression. For instance, Yang et al. [11] innovatively introduced a conditional diffusion model as decoder for image compression. Kuang et al. [26] proposed a consistency guidance architecture to guide the diffusion model in stably reconstructing high-quality images.
II-B Extreme Image Compression
@@
In recent years, extreme image compression has garnered increasing attention, aiming to compress image to extremely low bitrates, often below 0.1 bpp, while maintaining visually acceptable image quality. Gao et al. [27] leveraged the information-lossless property of invertible neural networks to mitigate the significant information loss in extreme image compression. Jiang et al. [28] treated text descriptions as prior to ensure semantic consistency between the reconstructions and the raw images. In [29], Wei et al. achieved extreme image compression by rescaling images using extreme scaling factors. Lu et al. [30] combined continuous and codebook-based discrete features to reconstruct high-quality images at extremely low bitrates.
Inspired by the great success of T2I diffusion models in various image restoration tasks [31, 32], some methods have incorporated T2I diffusion models into extreme image compression frameworks. Lei et al. [14] utilized a pre-trained ControlNet [33] to reconstruct images based on corresponding short text prompts and binary contour sketches. Careil et al. [15] conditioned iterative diffusion models on vector-quantized latent image representations and textual image descriptions. In our prior work [16], we combined compressive VAEs with pre-trained T2I diffusion models to achieve realistic reconstructions at extremely low bitrates.
However, constrained by the inherent characteristics of diffusion models, these diffusion-based extreme image compression methods are limited in both fidelity and efficiency. In this paper, we propose a solution to these limitations through a relay residual diffusion framework and a fixed-step fine-tuning strategy.
II-C Relay Diffusion
@@
Conventional diffusion models, such as denoising diffusion probabilistic models (DDPM) [13] and its variants, operate by progressively adding and removing Gaussian noise through a multi-step Markov process. These models have achieved remarkable results in low-resolution scenarios but face substantial challenges in terms of computational efficiency and performance when applied to higher resolutions. To overcome this, cascaded diffusion methods [34, 35] decompose the image generation into multiple stages, with each stage responsible for super-resolution conditioning on the previous one. However, these methods still require complete resampling at each stage, leading to inefficiencies and potential mismatches among different resolutions.
Relay diffusion, as proposed by Teng et al. [36], extends the cascaded framework by continuing the diffusion process directly from the low-resolution output rather than restarting from pure noise, which allows the higher-resolution stages to correct artifacts from earlier stages. This design is particularly well-suited for tasks such as image restoration and image compression, where degraded images or features are available. For instance, PASD [37] and SeeSR [38] directly embed the LR latent into the initial random noise during the inference process to alleviate the inconsistency between training and inference. ResShift [17] further constructs a Markov chain that transfers between degraded and target features by shifting the residual between them, substantially improving the transition efficiency. However, its redesigned diffusion equation and noise schedule prevent it from leveraging the robust generative capability of pre-trained stable diffusion.
In this work, we directly derive a new relay residual diffusion equation from stable diffusion’s diffusion equation, effectively combining the efficiency of residual diffusion with the powerful generative capability of stable diffusion.
III Methodology
@@
Figure 2: The proposed RDEIC. We first map a raw image $\boldsymbol{x}$ into the latent space using the encoder $\mathcal{E}$ and then perform end-to-end lossy compression to get compressed latent features $\boldsymbol{z}_{c}$. We then use $\boldsymbol{z}_{c}$ with added noise as the starting point and apply a denoising process to reconstruct the noise-free latent feature $\boldsymbol{z}_{0}$. The decoder $\mathcal{D}$ maps $\boldsymbol{z}_{0}$ back to the pixel space, to get the reconstructed image $\hat{\boldsymbol{x}}$. (a) Vanilla diffusion framework that starts from pure noise. (b) The proposed relay residual diffusion framework that starts from compressed latent features with added noise.
III-A Overall Framework
@@
Fig. 2 shows an overview of the proposed RDEIC network. We first use an encoder $\mathcal{E}$ and analysis transform $g_{a}$ to convert the input image $\boldsymbol{x}$ to its latent representation $\boldsymbol{y}$. Then we perform hyper transform coding on $\boldsymbol{y}$ with the categorical hyper model [39] and use the space-channel context model $\mathcal{C}_{m}$ to predict the entropy parameters $(\boldsymbol{\mu},\boldsymbol{\sigma})$ to estimate the distribution of quantized latent representation $\boldsymbol{\hat{y}}$ [20]. The side information $\boldsymbol{l}_{p}$ is quantized through vector-quantization, i.e., $\hat{\boldsymbol{l}}_{p}$ is the mapping of $\boldsymbol{l}_{p}$ to its closest codebook entry. Subsequently, the synthesis transform $g_{s}$ is used to obtain the image content dependent features $\boldsymbol{z}_{c}$. Random noise is then added to $\boldsymbol{z}_{c}$, which is the starting point for reconstructing the noise-free latent features $\boldsymbol{z}_{0}$ through an iterative denoising process. The denoising process is implemented by a frozen pre-trained noise estimator $\epsilon_{sd}$ of stable diffusion with trainable control network for intermediate feature modulation. Finally, the reconstructed image $\boldsymbol{\hat{x}}$ is decoded from $\boldsymbol{z}_{0}$ using the decoder $\mathcal{D}$.
III-B Relay Residual Diffusion
@@
Following stable diffusion, existing diffusion-based extreme image compression methods obtain the noisy latent by adding Gaussian noise with variance $\beta_{t}\in(0,1)$ to the noise-free latent features $\boldsymbol{z}_{0}$:
$\boldsymbol{z}_{t}=\sqrt{\bar{\alpha}_{t}}\boldsymbol{z}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{t},\ t=1,2,\cdots,T,$ (1)
@@
where $\epsilon_{t}\sim\mathcal{N}(0,\boldsymbol{I})$, $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$. When $t$ is large enough, the noisy latent $\boldsymbol{z}_{t}$ is nearly a standard Gaussian distribution. In practice, $T$ is typically very large, e.g., 1000, and pure noise is set as the starting point for the reverse diffusion process. However, this approach is not optimal for the image compression task, where the compressed latent features $\boldsymbol{z}_{c}$ are available.
@@
To this end, we set the starting point to $\boldsymbol{z}_{N}=\sqrt{\bar{\alpha}_{N}}\boldsymbol{z}_{c}+\sqrt{1-\bar{\alpha}_{N}}\epsilon_{N}$, where $N\ll T$. Our relay residual diffusion is thus defined as:
$\boldsymbol{z}_{n}=\sqrt{\bar{\alpha}_{n}}(\boldsymbol{z}_{0}+\eta_{n}\boldsymbol{e})+\sqrt{1-\bar{\alpha}_{n}}\epsilon_{n},\ n=1,2,\cdots,N,$ (2)
where $\boldsymbol{e}$ denotes the residual between $\boldsymbol{z}_{c}$ and $\boldsymbol{z}_{0}$, i.e., $\boldsymbol{e}=\boldsymbol{z}_{c}-\boldsymbol{z}_{0}$, and $\{\eta_{n}\}_{n=1}^{N}$ is a weight sequence that satisfies $\eta_{1}\to 0$ and $\eta_{N}=1$. Moreover, according to Eq. (2), $\boldsymbol{z}_{n-1}$ can be sampled as:
$\boldsymbol{z}_{n-1}$ $=\sqrt{\bar{\alpha}_{n-1}}(\boldsymbol{z}_{0}+\eta_{n-1}\boldsymbol{e})+\sqrt{1-\bar{\alpha}_{n-1}}\epsilon_{n-1},$ (3)
$=\sqrt{\bar{\alpha}_{n-1}}\boldsymbol{z}_{0}+\sqrt{\bar{\alpha}_{n-1}}\eta_{n-1}\boldsymbol{e}+\underbrace{\sqrt{1-\bar{\alpha}_{n-1}}\epsilon_{n-1}}_{\sim\mathcal{N}(0,(1-\bar{\alpha}_{n-1})\boldsymbol{I})}.$ (4)
@@
Since the residual $\boldsymbol{e}$ is unavailable during inference, we refer to DDIM [40] and assume that $\boldsymbol{z}_{n-1}$ is a linear combination of $\boldsymbol{z}_{n}$ and $\boldsymbol{z}_{0}$:
$\boldsymbol{z}_{n-1}=k_{n}\boldsymbol{z}_{0}+m_{n}\boldsymbol{z}_{n}+\sigma_{n}\epsilon.$ (5)
Substituting $\boldsymbol{z}_{n}$, as defined in Eq. (2), into Eq. (5), we obtain:
$\begin{split}\boldsymbol{z}_{n-1}=(k_{n}+m_{n}\sqrt{\bar{\alpha}_{n}})\boldsymbol{z}_{0}+m_{n}\sqrt{\bar{\alpha}_{n}}\eta_{n}\boldsymbol{e}+\\ \underbrace{m_{n}\sqrt{1-\bar{\alpha}_{n}}\epsilon_{n}+\sigma_{n}\epsilon}_{\sim\mathcal{N}(0,(m_{n}^{2}(1-\bar{\alpha}_{n})+\sigma_{n}^{2})\boldsymbol{I})}.\end{split}$ (6)
By combining Eq. (4) and Eq. (6), we obtain the following equations:
Note that, referring to DDIM [40], we set $\sigma_{n}=0$ for simplicity. By solving Eq. (7), we have: